OpenStudy (anonymous):

For the given supply and demand equations, find the consumer surplus. Demand Supply P=100 - 0.05x P=25 + 0.1 x

OpenStudy (anonymous):

p= 100 -0.05x multiplying with 20 on both side , we get 20p=2000-x

OpenStudy (anonymous):

Why do you multiply with 20?

OpenStudy (anonymous):

to make the equation easier

OpenStudy (anonymous):

Ok I guess I don't understand were the 20 came from I thought you were suppossed to set the equations equal to eachother....

OpenStudy (anonymous):

then for the supply equation , we have to multiply it with 10 on both sides which wud be 10p=250+x

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

solve for p and x p=75 and x=500

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

because we are applyling tjhe equilibrium condition

OpenStudy (anonymous):

because we are applyling tjhe equilibrium condition

OpenStudy (anonymous):

so I got the p=75 and x=500 a different way but where do i go from there

OpenStudy (anonymous):

now we find the consumer surplus triangle area where height would be 100 -75

OpenStudy (anonymous):

25

OpenStudy (anonymous):

base would be quantity traded which is 25 from from the supply equation

OpenStudy (anonymous):

so 0.5*25*25=312.5

OpenStudy (anonymous):

uh ok so none of my answers that I can chose from match up with that...?

OpenStudy (anonymous):

I think this is the method.....sorry if its wrong

OpenStudy (anonymous):

A.$6,250 B.$12,500 C.$19,000 D.$25,000 E.$26,000

OpenStudy (anonymous):

Those are the answers I can chose from....

OpenStudy (anonymous):

lemme c it once again then

OpenStudy (anonymous):

For the given supply and demand equations, find the consumer surplus. Demand P=100 - 0.05x Supply P=25 + 0.1 x

OpenStudy (anonymous):

sorry it is 6250

OpenStudy (anonymous):

the base is not 25 its 500 right which is the quanity traded

OpenStudy (anonymous):

so 500*25*0.5=6250

OpenStudy (anonymous):

okay with that

OpenStudy (anonymous):

Okay that works thank you!